OpenStudy (ashley_f97):

Differentiate the function: y=ln(lnx^3)

OpenStudy (superdavesuper):

do u know differentiation by substitution? u need that to solve this prob.

OpenStudy (ashley_f97):

Can you guide me through it?

OpenStudy (superdavesuper):

ok put z=lnx^3, then y=lnz right?

OpenStudy (ashley_f97):

Where did you get z?

OpenStudy (superdavesuper):

i just pick it because z=lnx^3 will make the differentiation simpler :)

OpenStudy (ashley_f97):

okay

OpenStudy (superdavesuper):

what is dy/dz then?

OpenStudy (ashley_f97):

Do you want me to actually put the dy/dz?

OpenStudy (ashley_f97):

Wait never mind. dz=3/x and dy=1/x

OpenStudy (superdavesuper):

Hahaha smart lady :P Yes I do want to calculate dy/dz because.... in the next step, we will use differentiation by substitution so that dy/dx=dy/dz * dz/dx

OpenStudy (superdavesuper):

wait y=lnz so dy/dz=?

OpenStudy (ashley_f97):

1/x/3/x?

OpenStudy (superdavesuper):

i think u are jumping ahead here.... y=lnz so dy/dz=1/z okay?

OpenStudy (ashley_f97):

okay

OpenStudy (superdavesuper):

Then like I said by substitution, dy/dx=(dy/dz)*(dz/dx) ok?

OpenStudy (ashley_f97):

ok

OpenStudy (superdavesuper):

so we already have dy/dz=1/z=1/lnx^3=1/(3*lnx) so we need to calculate dz/dx okay?

OpenStudy (ashley_f97):

okay

OpenStudy (superdavesuper):

z=ln(x^3) what is dz/dx?